During the constant volume reversible non-flow process with air, the temperature drops from 400K to 320K. Show process on the p-V and T-s coordinates and find the heat transfer in kJ/kg.
1
Expert's answer
2020-05-14T14:37:36-0400

Here, "T_1= 400 K, T_2= 320 K"

Heat flow at constant process can be find as

"H= n C_vdT = 1\times 0.718\times ( 400-320)= 57.44 J"